OpenStudy (anonymous):

20x to the 2nd power y to the 4th power over 5x to the 2nd pover y to the 3rd power

OpenStudy (anonymous):

= 4y

OpenStudy (anonymous):

20(x^2)(y^4)/5(x^2)(y^3)--> here we notice that 20/5 is 4 and x^2/x^2 using exponent laws we have 2-2=0 cancel out and (y^4)/(y^3) using exponent laws we notice that its 4-3=1 therefore we have just y^1=y